# From Sparse Dependence to Sparse Attention: Unveiling How Chain-of-Thought Enhances Transformer Sample Efficiency

### v1

@@ A.4.4 Third Step: Moving MLP in Linear Regime
$\left|\left\langle\frac{d\mathcal{L}^{(2)}}{dW_{r,1:d}}\right\rangle\right|\leq$ $\frac{\sqrt{\log(300nB/\delta)}}{m\sqrt{B}}.$
$\frac{d\mathcal{L}^{(2)}}{dW_{r,d+1:2d}}=$ $h_{r}\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\bm{1}(\nu_{r,i,b_{2}}>0)(-1)^{b_{1}+b_{2}}e_{S[i],b_{1}}+O(\frac{1}{n^{8}}).$
@@
$\left|\left\langle\frac{d\mathcal{L}^{(2)}}{dW_{r,d+1:2d}},e_{S[i^{\prime}],b^{\prime}}\right\rangle-\mathrm{sign}(\nu_{r,i,b^{\prime}})h_{r}\bm{1}(\nu_{r,i,0}\nu_{r,i,1}<0)\right|\leq\frac{1}{200m}$
@@
$\left|\left\langle W_{r,d+1:2d}^{(2)},e_{S[i^{\prime}],b^{\prime}}\right\rangle+\mathrm{sign}(\nu_{r,i,b^{\prime}})h_{r}\eta_{2}\bm{1}(\nu_{r,i,0}\nu_{r,i,1}<0)\right|<\frac{(\eta_{0}+\eta_{1})}{20mn}+\frac{\eta_{2}}{200m}<\frac{\eta_{2}}{100m}.$
@@ A.5 Extension to Hinge Loss
A.6 Technical Lemma
@@
Define $v=\sum_{i=1}^{n+k}\sum_{b=0}^{1}\lambda_{i,b}e_{i,b}$, then with probability $1-\delta$,
@@
Appendix B Additional Experiment Results
@@ B.1 Experiment Details
In this section, we analyze the normalized attention entropy of Qwen2-7B [68] and Qwen2-Math-7B models [49] on the GSM8K dataset [12] with and without CoT prompting respectively. Both models consist of $28$ layers, each with $28$ attention heads. The normalized attention entropy is computed as the average across the GSM8K test set. While Figure 6 presents the normalized attention entropy for specific attention heads in the first, last, and two intermediate layers, Figure 7 shows the entropy across all layers.
@@ B.2 Additional Results
In Section 4.2, we observe that training with repeated data can help Transformers learn the parity function, but it still requires significantly more computation compared to trained on CoT data (Figure 4). To further substantiate this observation, we explore the training of models with and without CoT on the $(n=20,k=12)$ parity problem. Compared with the $(n=20,k=6)$ parity problem considered in 4, this problem is harder for the models to learn as the number of secret variables $k$ is larger. We examine various configurations: the number of layers and heads ranges from $1,2,3,4,6$, and $8$; the learning rate varies from $6\times 10^{-6}$, $8\times 10^{-8}$, to $1\times 10^{-4}$; and the training dataset size varies from $10^{4}$, $10^{5}$ to $10^{6}$, with corresponding epochs ranging from $1000$, $100$ to $10$. Across all configurations we examined, training without CoT fails to achieve non-trivial accuracy (Figure 9). In contrast, models trained with CoT achieve perfect evaluation accuracy when trained on a dataset with $10,000$ samples for only $6$ epochs, or with approximately $60,000$ fresh samples in one-pass training setting.

### v2

@@ B.4.4 Third Step: Moving MLP in Linear Regime
$\frac{d\mathcal{L}^{(2)}}{dW_{r,d+1:2d}}=$ $h_{r}\sum_{i=n+1}^{n+k}\sum_{b_{1}=0}^{1}\sum_{b_{2}=0}^{1}\frac{1}{i}\bm{1}(\nu_{r,i,b_{2}}>0)(-1)^{b_{1}+b_{2}}e_{S[i],b_{1}}+O(\frac{1}{n^{8}}).$
@@
$\left|\left\langle\frac{d\mathcal{L}^{(2)}}{dW_{r,d+1:2d}},e_{S[i^{\prime}],b^{\prime}}\right\rangle-\mathrm{sign}(\nu_{r,i^{\prime},b^{\prime}})h_{r}\bm{1}(\nu_{r,i^{\prime},0}\nu_{r,i^{\prime},1}<0)\right|\leq\frac{1}{200m}$
@@
$\left|\left\langle W_{r,d+1:2d}^{(2)},e_{S[i^{\prime}],b^{\prime}}\right\rangle+\mathrm{sign}(\nu_{r,i^{\prime},b^{\prime}})h_{r}\eta_{2}\bm{1}(\nu_{r,i^{\prime},0}\nu_{r,i^{\prime},1}<0)\right|<\frac{(\eta_{0}+\eta_{1})}{20mn}+\frac{\eta_{2}}{200m}<\frac{\eta_{2}}{100m}.$
@@ B.5 Final Proof
Proof of Theorem 3.
The original theorem is a combination of Lemmas 27, 26 and 18. ∎
B.6 Technical Lemma
@@
Define $v=\sum_{i=n+1}^{n+k}\sum_{b=0}^{1}\lambda_{i,b}e_{i,b}$, then with probability $1-\delta$,
@@
Lemma 32.
If for constant $C$, the attention score before softmax has the following property:
  • for each position $i$, the target position $j=S[i]$ satisfies:
    $\left\langle e_{j,b_{1}^{\prime}},A^{(2)}e_{i,b_{2}^{\prime}}\right\rangle\geq 2C\log n.$
  • For all other positions $j^{\prime}\neq j$, the attention weights satisfy:
    $\left|\left\langle e_{j^{\prime},b_{1}^{\prime}},A^{(2)}e_{i,b_{2}^{\prime}}\right\rangle\right|\leq C\log n.$
The attention output is approximately one-hot, with
$\left|\mathcal{A}[A^{2}](\mathbf{E}(\textbf{b}))[i]-e_{S[i],\textbf{b}[S[i]]}\right|<4/n^{C-1}.$
Proof of Lemma 22.
The attention output for position $i$ is given by:
$\mathcal{A}[A^{2}](\mathbf{E}(\textbf{b}))[i]=\mathrm{X}\cdot\mathrm{softmax}\left(\mathrm{X}^{\top}A\mathrm{X}\right)[i],$
where the softmax is applied column-wise.
Given the condition from Lemma 5, we set
$\Delta=2C\log n$
This implies that:
$e^{-\Delta/2}<e^{-C\log n}=\frac{1}{n^{C}}.$
Define $z_{j^{\prime}}=\left\langle e_{j^{\prime},b_{1}^{\prime}},A^{(2)}e_{i,b_{2}^{\prime}}\right\rangle$.
The softmax for the target position $j=S[i]$ is:
$\mathrm{softmax}(z)_{j}=\frac{e^{z_{j}}}{e^{z_{j}}+\sum_{j^{\prime}\neq j}e^{z_{j^{\prime}}}}.$
Given that $z_{j}\geq\Delta$ and $|z_{j^{\prime}}|\leq\frac{\Delta}{2}$ (since $\frac{\eta_{0}\eta_{1}}{512mn^{2}}=\frac{\Delta}{2}$), we have:
$z_{j}-z_{j^{\prime}}\geq\Delta-\frac{\Delta}{2}=\frac{\Delta}{2}.$
Thus:
$\mathrm{softmax}(z)_{j}\geq\frac{e^{\Delta}}{e^{\Delta}+(T-1)e^{\Delta/2}}=\frac{1}{1+(T-1)e^{-\Delta/2}}.$
Therefore with $T\leq 2n$:
$\mathrm{softmax}(z)_{j}\geq\frac{1}{1+(T-1)\cdot\frac{1}{n^{C}}}\geq 1-\frac{2}{n^{C-1}}.$
Therefore
$\sum_{j^{\prime}\neq j}\mathrm{softmax}(z)_{j^{\prime}}\leq\frac{2}{n^{C-1}}.$
The attention output for position $i$ is:
$\mathcal{A}[A^{2}](\mathbf{E}(\textbf{b}))[i]=\sum_{j}e_{j,\textbf{b}[j]}\mathrm{softmax}(z)_{j}.$
Substituting the bounds:
$\left|\mathcal{A}[A^{2}](\mathbf{E}(\textbf{b}))[i]-e_{S[i],\textbf{b}[S[i]]}\right|$ $=\left|\sum_{j^{\prime}\neq S[i]}e_{j^{\prime},\textbf{b}[j^{\prime}]}\mathrm{softmax}(z)_{j^{\prime}}\right|+\left|e_{S[i],\textbf{b}[S[i]]}(\mathrm{softmax}(z)_{S[i]}-1)\right|$
$\leq\sum_{j^{\prime}\neq S[i]}\mathrm{softmax}(z)_{j^{\prime}}+\left|(\mathrm{softmax}(z)_{S[i]}-1)\right|\leq\frac{4}{n^{C-1}}$
We conclude our proof. ∎
Appendix C Additional Experiment Results
@@ C.1 Experiment Details
This section examines the normalized attention entropy of Qwen2-7B [71] and Qwen2-Math-7B models [52] on the GSM8K dataset [13] with and without CoT prompting respectively. Both models consist of $28$ layers, each with $28$ attention heads. The normalized attention entropy is computed as the average across the GSM8K test set. In Figure 6, the entropy values of each attention head in a layer are sorted separately under the three setups. While Figure 6 presents the normalized attention entropy for specific attention heads in the first, last, and two intermediate layers, Figure 7 shows the entropy across all layers.
@@ C.2 Additional Results
In Section 4.2, we observe that training with repeated data can help Transformers learn the parity function, but it still requires significantly more computation compared to trained on CoT data (Figure 4). To further substantiate this observation, we explore the training of models with and without CoT on the $(n=20,k=12)$ parity problem. Compared with the $(n=20,k=6)$ parity problem considered in Figure 4, this problem is harder for the models to learn as the number of secret variables $k$ is larger. We examine various configurations: the number of layers and heads ranges from $1,2,3,4,6$, and $8$; the learning rate varies from $6\times 10^{-6}$, $8\times 10^{-8}$, to $1\times 10^{-4}$; and the training dataset size varies from $10^{4}$, $10^{5}$ to $10^{6}$, with corresponding epochs ranging from $1000$, $100$ to $10$. Across all configurations we examined, training without CoT fails to achieve non-trivial accuracy (Figure 9). In contrast, models trained with CoT achieve perfect evaluation accuracy when trained on a dataset with $10,000$ samples for only $6$ epochs, or with approximately $60,000$ fresh samples in one-pass training setting.